two numbers total 71 & have a difference of 17, find the 2 numbers
Answer by rwm(914)

x+y=71
x-y=17
add the two
2x=88
x=44
y=71-44=27
y=27
another way to look at
the two numbers are the same except we add 17 to one of them
2a+17=71
2a=54
a=27
add 17 to it
b=44